OpenStudy (anonymous):

Help!!! Write an equation that you could solve using these steps: Add 7 to both sides. Then, divide both sides by 4. Explain how you know your answer is right.

OpenStudy (anonymous):

in words plz if u can

OpenStudy (anonymous):

2=1+1 5+7=1+4+7 5+7/4=1+4+7/4 12/4=12/4 3=3

OpenStudy (anonymous):

I am not 100% sure what you are asking for, but this is what I thought.

OpenStudy (anonymous):

u thought right OMG thank you so much

OpenStudy (anonymous):

sorry the top equation should be 5=4+1 not 2=1+1

OpenStudy (anonymous):

glad I could help